Two manned satellites approaching one another at a relative speed of 0.300 m/s intend to dock. The first has a mass of 3.00 ✕ 103 kg, and the second a mass of 7.50 ✕ 103 kg. Assume that the positive direction is directed from the second satellite towards the first satellite.

(a) Calculate the final velocity after docking, in the frame of reference in which the first satellite was originally at rest.

(b) What is the loss of kinetic energy in this inelastic collision?

(c) Repeat both parts, in the frame of reference in which the second satellite was originally at rest. Explain why the change in velocity is different in the two frames, whereas the change in kinetic energy is the same in both. (Do this on paper. Your instructor may ask you to turn in this work.)
Answer #1

(a) m2*v2 + m1*0 = (m1+m2)*v
7500kg * 0.300m/s + 0 = 10500kg * v
v = 0.214 m/s

(b) KEi = ½ * 7500kg * (0.300m/s)² = 337.5 J
KEf = ½ * 10500kg * (0.214m/s)² = 240.429 J
so KE loss = 97.071 J

(c) 0 - 3000kg * 0.300m/s = 10500kg * v
v = -0.0857 m/s

KEi = ½ * 3000kg * (-0.300m/s)² = 135 J
KEf = ½ * 10500kg * (0.0857/s)² = 38.558 J
So the KE loss is 96.442 J